OpenStudy (anonymous):

I manage a drug-testing program for an organization with 400 employees. Every three months, a random-number generator selects 100 names for testing. Afterward, these names go back into the selection pool. Obviously, the probability of an employee being chosen in one quarter is 25 percent. But what’s the likelihood of being chosen over the course of a year?

OpenStudy (anonymous):

this can be done in under 20 lines of Python!

OpenStudy (anonymous):

being chosen at least 1 time?

OpenStudy (anonymous):

\[p=\frac{1}{4}\] \[n=4\] \[P(X\geq1)=1-P(X=0)=1-\left(\frac{3}{4}\right)^4\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

how many times you will randomize it to get exact answer?

OpenStudy (anonymous):

in python

OpenStudy (wasiqss):

i will try it in c